### v1

# Cartesian Forest Matching

Abstract
In this paper, we introduce the notion of Cartesian Forest, which generalizes Cartesian Trees, in order to deal with partially ordered sequences. We show that algorithms that solve both exact and approximate Cartesian Tree Matching can be adapted to solve Cartesian Forest Matching in average linear time. We adapt the notion of Cartesian Tree Signature to Cartesian Forests and show how filters can be used to experimentally improve the algorithm for the exact matching. We also show a one to one correspondence between Cartesian Forests and Schröder Trees.
Keywords:
Cartesian Tree Cartesian Forest Pattern Matching Approximate Pattern Matching Schröder Tree
1 Introduction
Pattern matching consists of searching for one or all the occurrences of a pattern in a text. It is an essential task in many computer science applications. It can take different forms. For instance it can be done online when the pattern can be preprocessed or offline when the text can be preprocessed. Occurrences can be exact or approximate. When the pattern and the text are sequences of characters, it is known as string matching. When searching patterns in time series data, the notion of pattern matching is a bit more involved. Solutions can use Cartesian Trees that where introduced by Vuillemin in 1980 [19].
Cartesian Tree Matching has been introduced by Park et al. [15, 16]. Given a pattern, it consists of finding the factors of a text that share the same Cartesian Tree as the Cartesian Tree of the pattern.
Since then it has gained a lot of interest. Efficient solutions for practical cases for online search were given in [17]. Expected linear time algorithms are given in [1] for approximate Cartesian Tree Matching with one difference.
Indexing structures in the Cartesian Tree pattern matching framework are presented in [12, 10, 14]. Methods for computing regularities are given in [9] and methods for computing palindromic structures are presented in [6]. An algorithm for episode matching (given two sequences $p$ and $t$, finding all minimal length factors of $t$ that contains $p$ as a subsequence) in Cartesian Tree framework is presented in [13]. Practical methods for finding longest common Cartesian substrings of two strings appeared in [5]. Very recently, dynamic programming approaches for approximate Cartesian Tree pattern matching with edit distance has been considered in [11] and longest common Cartesian Tree subsequences are computed in [18]. Efficient algorithms for determining if two equal-length indeterminate strings match in the Cartesian Tree framework are given in [8].
Cartesian Trees are defined on arbitrary sequences where there can be values with multiple occurrences. In that case, ties can be broken in any way. In this article, in order to better deal with equal values we introduce Cartesian Forests.
We show that algorithms that solve both exact and approximate Cartesian Tree Matching can be adapted to solve Cartesian Forest Matching in average linear time. We then show a one to one correspondence between Cartesian Forests and Schröder Trees. More specifically we give algorithms for computing a Schröder Tree and a special type of words with parentheses given a Cartesian Forest. We also adapt the notion of Cartesian Tree Signature introduced in [4] to Cartesian Forests, and show how this notion can be used to experimentally improve the computation.
2 Definitions
In this paper, a sequence is always defined on an ordered alphabet. For a given sequence $x$, $|x|$ denotes the length of $x$. A sequence $v$ is factor of a sequence $x$ if $x=uvw$ for any sequences $u$ and $w$. A sequence $u$ is a prefix (resp. suffix) of a sequence $x$ if $x=uv$ (resp. $x=vu$). For a sequence $x$ of length $m$, $x[i]$ is the $i$-th element of $x$ and $x[i\ldots j]$ represents the factor of $x$ starting at the $i$-th element and ending at the $j$-th element, for $1\leq i\leq j\leq m$.
2.1 Cartesian Tree and Cartesian Forest
Definition 1 (Cartesian Tree $C(x)$)
Given a sequence $x$ of length $m$, the Cartesian Tree of $x$, denoted by $C(x)$, is the binary tree recursively defined as follows:
  • if $x$ is empty, then $C(x)$ is the empty tree;
  • if $x[1\ldots m]$ is not empty and $x[h]$ is the smallest value of $x$, $C(x)$ is the binary tree with $h$ as its root, the Cartesian Tree of $x[1\ldots h-1]$ as the left subtree and the Cartesian Tree of $x[h+1\ldots m]$ as the right subtree.
We denote by $\operatorname{rb}(T)$ the list of nodes on the right branch of a Cartesian Tree $T$. The Cartesian Tree of a sequence can be built online in linear time and space [7]. Informally, let $x$ be a sequence such that $C(x[1\ldots h-1])$ is already known. In order to build $C(x[1\ldots h])$, one only needs to find the nodes $j_{1}<\cdots<j_{k}$ in $\operatorname{rb}(C(x[1\ldots h-1]))$ such that $x[j_{1}]>x[h]$. If $j_{1}$ is the root of $C(x[1\ldots h-1])$ then $h$ becomes the root of $C(x[1\ldots h])$ otherwise let $j_{0}$ be the parent node of $j_{1}$, then node $h$ will be the root of the right subtree of $j_{0}$. In both cases, $j_{1}$ will be the root of the left subtree of node $h$. Then $\operatorname{rb}(C(x[1\ldots h])=\operatorname{rb}(C(x[1\ldots h-1]))\setminus(j_{1},\ldots,j_{k})\cup(h)$. All these operations can be easily done by implementing $\operatorname{rb}$ with a stack. The amortized cost of such an operation can be shown to be constant.
Cartesian Trees are defined on arbitrary sequences where there can be multiple occurrences of the smallest value.In that case, ties can be broken in any way. Usually the first occurrence of the smallest value is chosen to be the root of the tree. In order to better deal with equal values we now introduce Cartesian Forests. We first introduce the combinatorial object in itself, then we define what the Cartesian Forest of a sequence is.
Definition 2 (Cartesian Forest $F$)
A Cartesian Forest $F$ can be:
  • empty,
  • a sequence of $k$ planar trees rooted in $(r_{1},\ldots,r_{k})$, with $k\geq 1$, such that $r_{1}$ has a left Cartesian sub-Forest and a right Cartesian sub-Forest, and $r_{i}$ has only a right Cartesian sub-Forest for all $i\in\{2,\ldots,k\}$ (the left sub-Forest is necessarily empty).
Figure 1 shows examples of one Cartesian Forestand two forests that are not Cartesian.
Figure 1: On the left is a valid Cartesian Forest. In the middle, it is not a valid Cartesian Forest because the second tree in the right sub-Forest of $r^{\prime}_{1}$ has a left sub-Forest. On the right, a planar forest is not Cartesian: the second tree has a left sub-Forest.
Definition 3 (Cartesian Forest $F(x)$ of a sequence $x$)
Given a sequence $x$ of length $m$, the Cartesian Forest of $x$, denoted by $F(x)$, is recursively defined as follows:
  • if $x$ is empty, then $F(x)$ is the empty forest;
  • if $x[1\ldots m]$ is not empty, let $(r_{1},\ldots,r_{k})$ be the ordered sequence of all the $k$ positions of the smallest value of $x$, then $F(x)$ is a forest composed of $k$ planar trees whose roots are $r_{1},\ldots,r_{k}$, such that:
    • the Cartesian Forest $F(x[1\ldots r_{1}-1])$ forms a sequence of left subtrees of $r_{1}$,
    • for $1\leq i<k$, the Cartesian Forest $F(x[r_{i}+1\ldots r_{i+1}-1])$ forms a sequence of right subtrees of $r_{i}$,
    • the Cartesian Forest $F(x[r_{k}+1\ldots m])$ forms a sequence of right subtrees of $r_{k}$.
Figure 2 shows an example of the Cartesian Forest of a sequence.
Figure 2: Cartesian Forest associated to an ordered sequence $x$.
3 Exact Cartesian Forest Matching using Linear Representations
We adapt the linear representations of Cartesian Trees, such as the Parent-Distance [15] and the Skipped-Number representation [4] to Cartesian Forests. Basically, for a sequence $x$, when computing $C(x[1\ldots h])$ from $C(x[1\ldots h-1])$, the Parent-Distance of $h$ is equal to the distance between $h$ and its parent in $C(x[1\ldots h])$ and its Skipped-Number is the number of nodes removed from the right branch of $C(x[1\ldots h-1])$ comparing to the right branch of $C(x[1\ldots h])$ (see [1]).
The main idea of the adaptation to Cartesian Forests is simple: both representations involve the comparisons of two values $x[i]$ and $x[j]$ in the original sequence $x$, for a given $i<j$. Suppose we have either $x[i]>x[j]$ or $x[i]<x[j]$, then we are in the case of Cartesian Tree and the linear representation is the same.
The idea for the Parent-Distance of the Cartesian Forest of a sequence $x$ is that its absolute value at a position $h$ gives the distance between a node $h$ and its parent or left sibling in the Cartesian Forest associated to the sequence $x[1\ldots h]$.
Formally, let $\operatorname{small}_{x}(h)=\max_{1\leq j<h}\{j\ |\ x[j]<x[h]\}\cup\{0\}$ and $\operatorname{equal}_{x}(h)=\max_{1\leq j<h}\{j\ |\ x[j]=x[h]\}\cup\{0\}$.
Definition 4 (Parent-Distance representation $PD_{x}$)
Given a sequence $x[1\ldots m]$, the Parent-Distance representation of $x$ is an integer sequence $PD_{x}[1\ldots m]$ defined as follows:
$PD_{x}[h]=\begin{cases}h-\operatorname{small}_{x}(h)&\text{if }\operatorname{small}_{x}(h)>\operatorname{equal}_{x}(h)\cr-(h-\operatorname{equal}_{x}(h))&\text{if }\operatorname{small}_{x}(h)<\operatorname{equal}_{x}(h)\cr 0&\text{otherwise.}\end{cases}$
Definition 5 (The referent table $\operatorname{ref}_{x}$)
Given a sequence $x[1\ldots m]$, the referent table of $x$ is a sequence of sets $\operatorname{ref}_{x}[1\ldots m]$ such that
$\operatorname{ref}_{x}[h]=\begin{cases}\min_{h<j\leq m}\{j\ |\ x[j]\leq x[h]\}&\text{if such }j\text{ exists}\cr-1&\text{ otherwise.}\end{cases}$
Definition 6 (Skipped-Number representation $SN_{x}$)
Given a sequence $x[1\ldots m]$, the Skipped-number representationof $x$ is an integer sequence $SN_{x}[1\ldots m]$ such that
$SN_{x}[h]=\begin{cases}|\{j<h\mid\operatorname{ref}_{x}[j]=h\}|&\mbox{if }\operatorname{small}_{x}(h)>\operatorname{equal}_{x}(h)\cr-|\{j<h\mid\operatorname{ref}_{x}[j]=h\}|&\mbox{if }\operatorname{small}_{x}(h)<\operatorname{equal}_{x}(h)\cr\end{cases}$
Figure 3 shows examples of Parent-Distance representations, Skipped-number representations and referent tables of two Cartesian Forests.
Figure 3: Two sequences $x$ and $y$, and their associated Cartesian Forests $F(x)$ and $F(y)$ and their corresponding Parent-Distance representations, Skipped-number representation and referent tables. As one can see, $x$ is a prefix of $y$ and the forest $F(x)$ is transformed into a sequence of left subtrees in $F(y)$.
Given two sequences $x$ and $y$, we will denote $x\approx_{CF}y$ when the two sequences share the same Cartesian Forest.
Definition 7 (Cartesian Forest Matching (CFM))
Given two sequences $p[1\ldots m]$ and $t[1\ldots n]$, find every position $j$, with $11\leq j\leq n-m+$ 1, such that $t[j\ldots j+m-1]\approx_{CF}p[1\ldots m]$.
Example 1
Let $t=(5,7,3,6,3,7,2,8,2,4,3,3)$ and $p=(2,3,1,4,1,5)$ respectively be the text and the pattern. We have two occurrences of $p$ in $t$ as
$t[1\ldots 6]=(5,7,3,6,3,7)\approx_{CF}p\text{ and }t[5\ldots 10]=(3,7,2,8,2,4)\approx_{CF}p.$
Input : Two sequences $p$ and $t$ of length $m$ and $n$
Output : The number of positions $j$ such that $p$ is equivalent to $t[j\ldots j+m-1]$
1 $occ\leftarrow 0$
2 $x\leftarrow t[1\ldots m]$
3 $LN_{p},LN_{x}\leftarrow$ linear representations of $C(p)$ and $C(x)$
4 for $j\in\{1,\ldots,n-m+1\}$ do
5       if $LN_{p}=LN_{x}$ then
6             $occ\leftarrow occ+1$
7      $x\leftarrow t[j+1\ldots j+m]$
8       Update $LN_{x}$
9
return $occ$
Algorithm 1 metaAlgorithm($p$, $t$)
Proposition 1
Algorithm 1 solves the CFM problem.
Proof
In Algorithm 1, that solves the CFM problem, $LN_{x}$ is a linear representations of the Cartesian Forest of a sequence $x$. It can indifferently be its Parent-Distance table or its Skipped-number representation. Note that this algorithm already appears in [1].
In [1] the authors show that this algorithm (originally applied to solve the Cartesian Tree Matching problem) has a $\mathcal{O}(mn)$ worst-case time-complexity and a $\mathcal{O}(m)$ space-complexity. It is also proved that the average-case time-complexity is $\mathcal{O}(n)$ in several random models.
The same results hold for Cartesian Forest Matching (even if some average-case results require an adaptation of the proofs, the main ideas remain).
4 Approximate Cartesian Forest matching
The above approach for exact CFM can be extended to approximate CFM with one difference in a similar way as approximate CTM with one difference is done in [1].
Given a sequence $x$ of length $m$ and $i$ a position $1\leq i<m$, let $y$ be the sequence defined by swapping the elements in positions $i$ and $i+1$ in $x$. In [1], it is shown that there are at most $3$ mismatches between the Skipped-number representation of the Cartesian Tree of $x$ and the Skipped-number representation of the Cartesian Tree of $y$. The same result holds for the Skipped-number representation of the Cartesian Forest of $x$ and the Skipped-number representation of the Cartesian Forest of $y$ considering that two elements match if their absolute values are equal. Thus, based on the results in [1], the CFM with one swap of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
Approximate CTM with one mismatch, one insertion or one deletion is solved by comparing the Parent-Distance representations from left to right and the Parent-Distance representations from right to left of the Cartesian Trees of $x$ and $y$. The Parent-Distance representation from right to left of a Cartesian Forest can be defined in a similar way as for a Cartesian Tree. And then, approximate CFM with one mismatch, one insertion or one deletion can be solved in a similar way than approximate CTM with one mismatch, one insertion or one deletion. Thus, based on the results in [1], the CFM with one mismatch, one insertion or one deletion of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
5 Combinatorics of the Cartesian Forests
5.1 Recursive Definition
According to Definition 2, a Cartesian Forest $F$ can be either empty (denoted by $\emptyset$) or contains at least one node. This node has a left sub-forest and a right sub-forest, but it can also have siblings. A sibling $S$ is a particular Cartesian Forest that cannot have a left sub-forest. Therefore, we obtain the following recursive decomposition:
[…]
5.2 Generating Function
Let $F(z)=\sum_{n\geq 0}f_{n}z^{n}$ be the generating function of Cartesian Forests, where $f_{n}$ counts the number of Cartesian Forests with $n$ nodes and $S(z)$ be the generating function of siblings. From the previous recursive decomposition we have:
$\begin{cases}F(z)=z\cdot F^{2}(z)\cdot S(z)+1\\ S(z)=z\cdot F(z)\cdot S(z)+1\end{cases}$
from which we obtain that
$S(z)=\frac{1}{1-z\cdot F(z)}\text{ and }F(z)=1+\frac{z\cdot F^{2}(z)}{1-z\cdot F(z)}$
It can be shown from here that $F(z)=\frac{1+z+\sqrt{1-6z+z^{2}}}{4z}$, whose associated coefficients are known to be Schröder–Hipparchus numbers, also called super-Catalan numbers (Sequence $A001003$ on OEIS) enumerated by the following formula:
$f_{n}=\sum_{i=1}^{n}\frac{1}{n}\binom{n}{i}\binom{n}{i-1}2^{i-1}$
which, amongst other things, counts the number of ways of inserting parentheses into a sequence of $n+1$ symbols, where each pair of parentheses surrounds at least two symbols or parenthesized groups, and without any parentheses surrounding the entire sequence. In [2], the authors show that:
$f_{n}\sim\frac{\sqrt{3\sqrt{2}-4}}{4\sqrt{n^{3}\pi}}(3+2\sqrt{2})^{n}\sim 0.07(5.828)^{n}n^{-\frac{3}{2}}.$
In the following subsections, we describe bijections between Cartesian Forests and classical combinatorial objects counted by the Schröder-Hipparchus Number. Figure 5 shows the one to one correspondence for $n\in\{1,2,3\}$ between Cartesian Trees, Schröder Trees and Parentheses Words.
5.3 A bijection with Schröder Trees
Definition 8 (Schröder Tree $ST$)
A Schröder Tree is a planar tree whose internal nodes have two or more subtrees.
Schröder Trees with $n+1$ leaves are counted by $f_{n}$.
Given a Cartesian Forest $F$ with $k$ roots $(r_{1},\ldots,r_{k})$, we denote $\operatorname{left}(r_{1})$ the left sub-forest of the root $r_{1}$. For $i\in\{1,\ldots,k\}$, we denote $\operatorname{right}(r_{i})$ the right sub-forest of $r_{i}$.
From a Cartesian Forest with $n\geq 1$ nodes, Algorithm 2 gives a recursive approach to building a Schröder Tree with $n+1$ leaves .
Input : A Cartesian Forest $F$ with $n$ nodes and $k$ roots, s.t. $F=(r_{1},\ldots,r_{k})$
Output : A Schröder Tree $ST$ with $n+1$ leaves
1 $ST\leftarrow$ a root;
2 if $F$ is not empty then
3       $(c_{1},\ldots,c_{k+1})\leftarrow$ Create a tuple of subtrees;
4       $c_{1}\leftarrow$ CFtoST$(\operatorname{left}(r_{1}))$;
5       for $i\in\{2,\ldots,k+1\}$ do
6             $c_{i}\leftarrow$ CFtoST$(\operatorname{right}(r_{i-1}))$;
7
8      $ST\leftarrow$ Add subtrees $(c_{1},\ldots,c_{k+1})$ to $ST$;
9
10return $ST$;
Algorithm 2 CFtoST$(F)$
Input : A Cartesian Forest $F$ with $n$ nodes and $k$ roots, s.t. $F=(r_{1},\ldots,r_{k})$
Output : A sequence $w$ with parentheses
1 $w\leftarrow\square$;
2 if $F$ is not empty then
3       $(w_{1}\cdots w_{k+1})\leftarrow$ Create a tuple of words ;
4       $w_{1}\leftarrow\textsc{CFtoW}(\operatorname{left}(r_{1}))$;
5       for $i\in\{2,\ldots,k+1\}$ do
6             $w_{i}\leftarrow\textsc{CFtoW}(\operatorname{right}(r_{i-1}))$;
7
8      $w\leftarrow\bm{(}w_{1}\cdot w_{2}\cdots w_{k+1}\bm{)}$;
9
10return w;
Algorithm 3 CFtoW($F$)
Figure 4: Both algorithms take a Cartesian Forest as input and return the corresponding Schröder Tree (on the left) or Parentheses Word (on the right). As one can see, both methods are very similar.
Algorithm 2 is an injective function: at each level, the number of subtrees added to the Schröder Tree is exactly $k+1$, where $k$ is the number of trees in the Cartesian Forest. Since both objects share the same generating function, we obtain the result announced in Lemma 1.
5.4 A bijection with Parentheses Words
Definition 9 (Parentheses Word $w$)
A Parentheses Word $w$ is a word: over the alphabet {$\{\bm{(}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \square\leavevmode\nobreak\ ,\leavevmode\nobreak\ \bm{)}\}$ such that
  • either $w=\square$
  • or $w=\bm{(}w_{1}\cdots w_{k}\bm{)}$ where $k\geq 2$ and each $w_{i}$ is a parentheses word.
Do note that, in this definition, unlike the more commonly found definition of these words, we allow parentheses to surround the entire sequence. We notably do so in order to simplify Algorithm 3 and to make the bijection more apparent to the reader. But since every parentheses word that is not . is contained between a common parentheses, it is not necessary to represent it. Hence, in Figure 5, we omit to draw those parentheses in order to match classical representations (we write $\bm{(}\square\square\bm{)}\square$ instead of $\bm{(}\bm{(}\square\square\bm{)}\square\bm{)}$ for instance). Informally, one may consider the symbols of the word as separators between the nodes and each group of parentheses as a (sub-)forest.
Using the same arguments as for Schröder Tree, Algorithm 3 is an injective function.
Lemma 1
Algorithms 2 and 3 are bijective functions that map a Cartesian Forest with $n$ nodes to a Schröder Tree with $n+1$ leaves (Algorithm 2) or to a Parentheses Word with $n+1$ symbols (Algorithm 3). They both have a $\Theta(n)$ time-complexity and a $\Theta(n)$ worst-case space complexity.
Figure 5: Correspondence between Cartesian Forests with $n$ nodes, Schröder Trees with $n+1$ leaves and sequences of length $n+1$ with parentheses when $n\in\{1,2,3\}$.
6 Cartesian Forest Signature and Cartesian Forest Matching using a filter
In [4], the authors propose a signature, or perfect hash, of a Cartesian Tree, based on the Skipped-number representation. Given a Cartesian Tree with $n$ nodes, its signature is an integer with at most $2n$ bits. In this section, we extend this notion of signature to Cartesian Forests, obtaining a signature with at most $3n$ bits.
Definition 10 (Cartesian Forest Signature)
Given a Cartesian Forest $F$ and its Skipped-number representation $SN_{F}$ (see Definition 6), its signature is defined in the following way: given a position $i\in\{1,\ldots,m\}$ in the Skipped-number representation, each value $SN_{F}[i]$ is encoded by a sequence of bits, that are concatenated to obtain an integer.
  • The first bits concern the sign of $SN_{F}[i]$. If $SN_{F}[i]=0$, it is equal to 00. The case $SN_{F}[i]<0$ is encoded by $10$ and $SN_{F}[i]>0$ is encoded by $11$.
  • If $SN_{F}[i]\neq 0$, the following bits are a unary encoding of $|SN_{F}[i]|$: $|SN_{F}[i]|-1$ bits equal to $1$ followed by a bit equal to 00.
Since the total number of skipped nodes cannot exceed $m$, a signature contains at most $3m$ bits.
This representation could be used efficiently as a perfect hash for small patterns, to obtain an efficient algorithm to solve the Cartesian Forest Matching problem: one only needs to update the signature as one would update the Skipped-number representation of the chunk of text compared to the pattern, using bitwise operations. But it becomes inefficient when $3m$ exceeds the size of a register, as one cannot use those operations efficiently anymore.
Therefore, in order to accelerate the Cartesian Forest Matching in a Rabin-Karp fashion we now introduce a filtering method.
Definition 11 (Cartesian Forest $\tau$-Filter)
Given a sequence $x$ of length $m$ and its linear representation $LN_{x}$, its $\tau$-filter $\mathcal{F}il_{x}$ is a sequence of $\tau$ bits such that $\mathcal{F}il_{x}[i]=0$ if $LN_{x}[m-\tau+i]=0$ and $\mathcal{F}il_{x}[i]=1$ otherwise.
Algorithm 1 can be adapted: line $5$ is only tested if $\mathcal{F}il_{x}$ is equal to $\mathcal{F}il_{p}$. $\tau$ can be adapted to match the size of a classical register, that is $32$, $64$ or $128$ bits. Therefore, comparing two filters can be made in constant time. The update of the filter can also be made in constant time by tracking the positions in $LN_{x}$ that have been updated (see [1] for more details on the update function of the linear representations). In the following Section, we only implemented the filter for the Skipped-number representation.
7 Experiments
In this Section, we implemented a Cartesian $\tau$-filter using $\tau=64$. The first set of experiments consider the uniform distribution over text and patterns of respective length $n$ and $m$, over a $k$ letter alphabet. Figure 6 sums up the details and observations.
Figure 6: In this experiment, we randomly generated 10 0001000010\,00010 000 patterns of size $m$ and texts of size $1000$. From left to right, the alpĥabet is respectively equal to $2$, $4$ and finally $m$. As one, can see, the Parent-Distance version is the slowest, then comes the one using the Skipped-number representation. The method using a filter is the faster one, but only with a slight margin. The average cost slightly increases with $k$, which is probably due to the cost of the update function, since the average Parent-Distance increases with the size of the alphabet. On the contrary, the standard deviation seems to decreases, since we obtain smoother curves.
Efficient even with (not too) low entropy
Figure 7: Both figures represent the same experiment: the one on the left is the full one and the one on the right is a zoom on the results when the entropy is low (between $0.01$ and $0.3$). In this experiment, for each value of the collision entropy, 10 0001000010\,00010 000 random texts of length $1000$ and patterns of length $100$ were generated and the three algorithms were applied to them. The added value of using the filter method, compared to the simple Skipped-number representation version, is more important when the entropy is low, whereas the difference in efficiency between the Parent-Distance version and the Skipped-number representation version drops.
The following experiment uses the random generator from [3]. It generates sequences uniformly amongst those of a fixed length over a fixed alphabet and a fixed Collision Entropy (that is Rényi Entropy with $\alpha=2$). The Collision Entropy is a function of the probability for two random variables (a letter in the pattern $p$ and the text $t$) to be equal. The lower the entropy is, the higher average complexity of the algorithms. As a matter of fact, if the entropy reaches 00, then both the text and the pattern contains the repetition of a unique symbol, which corresponds to the worst-case complexity of Algorithm 1. Though, as one can see in Figure 7, the average cost of the algorithms quickly drops and stabilizes.