### v2

# Cartesian Forest Matching

Abstract
In this paper, we introduce the notion of Cartesian Forest, which generalizes Cartesian Trees, in order to deal with partially ordered sequences. We show that algorithms that solve both exact and approximate Cartesian Tree Matching can be adapted to solve Cartesian Forest Matching in linear time in the worst-case for exact matching and linear time on average for approximate matching. We adapt the notion of Cartesian Tree Signature to Cartesian Forests. We also show a one-to-one correspondence between Cartesian Forests and Schröder Trees.
1 Introduction
Pattern matching consists of searching for one or all the occurrences of a pattern in a text. It is an essential task in many computer science applications. It can take different forms. For instance it can be done online when the pattern can be preprocessed or offline when the text can be preprocessed. Occurrences can be exact or approximate. When the pattern and the text are sequences of characters, it is known as string matching. When searching patterns in time series data, the notion of pattern matching is a bit more involved. Solutions can use Cartesian Trees that where introduced by Vuillemin in 1980 [23].
Cartesian Tree Matching has been introduced by Park et al. [19, 20]. Given a pattern, it consists of finding the factors of a text that share the same Cartesian Tree as the Cartesian Tree of the pattern.
Since then it has gained a lot of interest. Efficient solutions for practical cases for online search were given in [21]. Expected linear time algorithms are given in [3] for approximate Cartesian Tree Matching with one difference.
Indexing structures in the Cartesian Tree pattern matching framework are presented in [16, 13, 18]. Methods for computing regularities are given in [12] and methods for computing palindromic structures are presented in [9]. An algorithm for episode matching (given two sequences $p$ and $t$, finding all minimal length factors of $t$ that contains $p$ as a subsequence) in Cartesian Tree framework is presented in [17]. Practical methods for finding longest common Cartesian substrings of two strings appeared in [7]. Very recently, dynamic programming approaches for approximate Cartesian Tree pattern matching with edit distance have been considered in [14] and longest common Cartesian Tree subsequences are computed in [22]. Efficient algorithms for determining if two equal-length indeterminate strings match in the Cartesian Tree framework are given in [11].
Cartesian Trees are defined (Section 2) on arbitrary sequences where there can be values with multiple occurrences. In that case, ties should be broken in some way. In this article, in order to better deal with equal values we introduce Cartesian Forests (Section 3).
We show that algorithms that solve both exact and approximate Cartesian Tree Matching (Section 4 and Section 5) can be adapted to solve Cartesian Forest Matching in linear time in the worst-case for exact matching and linear time on average for approximate matching. We then show a one to one correspondence between Cartesian Forests and Schröder Trees (Section 6). More specifically we give algorithms for computing a Schröder Tree and a special type of words with parentheses given a Cartesian Forest. We also adapt the notion of Cartesian Tree Signature introduced in [6] to Cartesian Forests (Section 7), and show how this notion can be used to experimentally improve the computation (Section 8).
2 Preliminaries
In this paper, a sequence is always defined on an ordered alphabet. For a given sequence $x$, $|x|$ denotes the length of $x$. A sequence $v$ is factor of a sequence $x$ if $x=uvw$ for any sequences $u$ and $w$. A sequence $u$ is a prefix (resp. suffix) of a sequence $x$ if $x=uv$ (resp. $x=vu$). A prefix of a sequence that is also a suffix is called a border. For a sequence $x$ of length $m$, $x[i]$ is the $i$-th element of $x$ and $x[i\ldots j]$ represents the factor of $x$ starting at the $i$-th element and ending at the $j$-th element, for $1\leq i\leq j\leq m$.
2.1 Cartesian Trees
Definition 1(Cartesian Tree $C(x)$)
Given a totally ordered sequence $x$ of length $m$, the Cartesian Tree of $x$, denoted by $C(x)$, is the binary tree recursively defined as follows:
  • if $x$ is empty, then $C(x)$ is the empty tree;
  • if $x[1\ldots m]$ is not empty and $x[i]$ is the smallest value of $x$, $C(x)$ is the binary tree with $i$ as its root, the Cartesian Tree of $x[1\ldots i-1]$ as the left subtree and the Cartesian Tree of $x[i+1\ldots m]$ as the right subtree.
We denote by $\operatorname{rb}(T)$ the list of nodes on the right branch of a Cartesian Tree $T$. Also, $C_{i}(x)=C(x[1\ldots i])$ to simplify the notations. The Cartesian Tree of a sequence can be built online in linear time and space [10]. Informally, let $x$ be a sequence such that $C_{i-1}(x)$ is already known. In order to build $C_{i}(x)$, one only needs to find the nodes $j_{1}<\cdots<j_{k}$ in $\operatorname{rb}(C_{i-1}(x))$ such that $x[j_{1}]>x[i]$. Then, $j_{1}$ will be the root of the left subtree of node $i$. If $j_{1}$ is the root of $C_{i-1}(x)$ then $i$ becomes the root of $C_{i}(x)$, otherwise let $j_{0}$ be the parent node of $j_{1}$, then node $i$ will be the root of the right subtree of $j_{0}$. Then $\operatorname{rb}(C_{i}(x))=\operatorname{rb}(C_{i-1}(x))\setminus(j_{1},\ldots,j_{k})\cup(i)$. All these operations can be easily done by implementing $\operatorname{rb}$ with a stack. The amortized cost of such an operation can be shown to be constant.
Cartesian Trees are defined on arbitrary sequences where there can be multiple occurrences of the smallest value. In that case, ties should be broken in some way. Usually the first occurrence of the smallest value is chosen to be the root of the tree.
Given a Cartesian Tree $C(x)$ of a sequence $x$ of length $m$, a linear representation of $C(x)$, denoted $\operatorname{LN}_{x}$, is a table of integers of length $m$, and the set of linear representations of a given length is in bijection with the set of Cartesian Trees of the same size. A linear representation of Cartesian Trees was introduced by Park et al. [19]: the parent-distance representation (see example Figure 3). Using this representation and classical string algorithms (Knuth-Morris-Pratt (KMP) [15] and Aho-Corasick [1] algorithms), they obtained linear-time solutions for single and multiple pattern Cartesian Tree matching (see Figure 1 for an example).
Let $\operatorname{small}_{x}(i)=\max(\{k\mid x[k]<x[i]\mbox{ with }1\leq k<i\}\cup\{0\})$.
Definition 2(Parent-distance representation $\operatorname{PD}_{x}$)
Given a sequence $x$ of length $m$, the parent-distance representation of $x$ is an integer sequence $\operatorname{PD}_{x}[1\ldots m]$, which is defined as follows:
$\operatorname{PD}_{x}[i]=\begin{cases}i-\operatorname{small}_{x}(i)&\text{if }\operatorname{small}_{x}(i)>0\cr 0&\text{otherwise.}\end{cases}$
Since the parent-distance representation has a one-to-one mapping with Cartesian Trees, it can replace them without loss of information.
The parent-distance table is not the only linear representation of Cartesian Trees. Another linear representation is based on the number of nodes skipped by a new node when building online the Cartesian Tree of a sequence. Informally, for each node $i$, it stores the number of nodes $(j_{1}<\cdots<j_{k})$ deleted from the right path when computing $C_{i}(x)$ from $C_{i-1}(x)$. We say that node $i$ skipped nodes $(j_{1}<\cdots<j_{k})$ and that these nodes are skipped by node $i$. The left subtree of a Cartesian Tree $C(x)$ is denoted $left(C(x))$ and the length of its right branch $\operatorname{RB}(C(x))=|\operatorname{rb}(C(x))|$.
See Figure 1 for an example.
Definition 3(Skipped-number representation $SN_{x}$)
Given a sequence $x[1\ldots m]$, the Skipped-number representation of $x$ is a sequence $SN_{x}[1\ldots m]$ of integers such that $SN_{x}[i]$ is the length of the right-branch of the left-subtree of the subtree rooted at node $i$ of the Cartesian Tree of $x$, that is $SN_{x}[i]=\operatorname{RB}(\operatorname{left}(C_{i}(x)))$.
Both linear representations can be computed in linear time [19, 6, 3].
2.2 Exact Cartesian Tree Matching
Given two sequences $x$ and $y$ of length $m$ and $n$ respectively, the Exact Cartesian Tree Matching problem (CTM) consists of finding all the factors of $y$ that have the same Cartesian Tree of the one of $x$. This can be done in time $O(m+n)$ by using a KMP-like algorithm. To achieve this complexity, a Cartesian border table is used. Given the Parent-Distance representation of $x$, the Parent-Distance of a factor $x[i\ldots j]$ of $x$ satisfies:
$\operatorname{PD}_{x[i\ldots j]}[k]=\begin{cases}0&\mbox{if }\operatorname{PD}_{x}[i+k-1]\geq k,\cr\operatorname{PD}_{x}[i+k-1]&\mbox{otherwise}.\end{cases}$
Definition 4(Cartesian Border Table $CTBord$)
The Cartesian border table of $x$ is defined in the following way: $CTBord[0]=0$ and, for $1\leq k<i$, $CTBord[i]=\max\{k\mid C_{k}(x)=C(x[i-k+1\ldots i])\}$.
Figure 1 shows the Cartesian border table of $x=(3,1,6,4,8,6,7,5,9$).
$i$ 1 2 3 4 5 6 7 8 9
$x[i]$ 3 1 6 4 8 6 7 5 9
$\operatorname{PD}_{x}[i]$ 0 0 1 2 1 2 1 4 1
$SN_{x}[i]$ 0 1 0 1 0 1 0 2 0
$CTBord[i]$ 0 1 1 2 3 4 5 2 3
Figure 1: The parent-distance representation, the skipped-number representation and the Cartesian border table of a sequence $x$.
2.3 Approximate Cartesian Tree Matching
Given two sequences $x$ and $y$ of length $m$ and $n$ respectively, the Approximate Cartesian Tree Matching problem (ACTM) consists of finding all the factors of $y$ that have the same Cartesian Tree of the one of $x$ up to some differences.
Unfortunately, the linear-time solutions obtained in [19] for the exact case do not work anymore in the general approximate case. Also we conjecture that there is no algorithm that solves approximate pattern matching with a linear worst-case complexity.
In [3], the authors provide the following MetaAlgorithm that can solve ACTM with up to one difference (substitution, insertion, deletion, swap) in quadratic worst-case time complexity and in linear time in average. The test at line $5$, using the $\sim$ symbol, signifies that depending on the considered notion of error, we will not test equality between linear representation but rather test an equivalence.
Input : Two sequences $p$ and $t$ of length $m$ and $n$
Output : The number of positions $j$ such that $p$ is equivalent to $t[j\ldots j+m-1]$
1 $occ\leftarrow 0$
2 $x\leftarrow t[1\ldots m]$
3 $\operatorname{LN}_{p},\operatorname{LN}_{x}\leftarrow$ linear representations of $C(p)$ and $C(x)$
4 for $j\in\{1,\ldots,n-m+1\}$ do
5 if $\operatorname{LN}_{p}\sim\operatorname{LN}_{x}$ then
6    $occ\leftarrow occ+1$
7 $x\leftarrow t[j+1\ldots j+m]$
8   Update $\operatorname{LN}_{x}$
9
return occocc
Algorithm 1 metaAlgorithm($p$, $t$)
3 Cartesian Forests
A limitation of Cartesian Trees is that it does not take into account equal values in the original sequence. For instance, the sequence $(1,1,1,1,1)$ shares the same Cartesian Tree as $(1,2,3,4,5)$ when ties are broken by ordering equal values according to their positions. In order to better deal with equal values we now introduce Cartesian Forests.
Recall that a planar tree, or general tree, is a tree in which each node possesses a sequence of children, or subtrees. In the following, trees designates planar trees.
Informally, during the construction of $C_{i}(x)$ from $C_{i-1}(x)$, assume that node $i$ should be inserted in the right branch between nodes $j_{0}$ and $j_{1}$ such that $x[j_{0}]\leq x[i]<x[j_{1}]$. In a Cartesian Tree, node $i$ becomes the root of the right subtree of node $j_{0}$ and node $j_{1}$ becomes the root of the left subtree of node $i$. In a Cartesian Forest, node $i$ becomes the root of the right subtree of node $j_{0}$ only if $x[j_{0}]<x[i]$ while it becomes the right sibling of node $j_{0}$ if $x[j_{0}]=x[i]$ and node $j_{1}$ remains in the right subtree of node $j_{0}$.
A forest is a sequence of trees. An empty forest is an empty sequence that does not contain any tree. The children of a node can be seen as a sub-forest. In the following, a node of a tree in a Cartesian Forest, instead of having a sequence of children, can possess a left sequence of children and a right sequence of children, which we call the left sub-forest and the right sub-forest. The leftmost tree of a Cartesian Forest is actually the only one that can possess a left sub-forest and a right sub-forest, whereas all the other trees can only have a right sub-forest. Figure 2 illustrates this with an example of a Cartesian Forest.
Definition 5(Cartesian Forest $F(x)$ of a sequence $x$)
Given a sequence $x$ of length $m$, the Cartesian Forest of $x$, denoted by $F(x)$, is recursively defined as follows:
  • if $x$ is empty, then $F(x)$ is the empty forest;
  • if $x[1\ldots m]$ is not empty, let $(r_{1},\ldots,r_{k})$ be the ordered sequence of all the $k$ positions of the smallest value of $x$ and let $r_{k+1}=m+1$, then $F(x)$ is a forest composed of $k$ trees whose roots are $r_{1},\ldots,r_{k}$, such that:
    • the left sub-forest of $r_{1}$ is the Cartesian Forest $F(x[1\ldots r_{1}-1])$, and
    • for $1\leq i\leq k$, the right sub-forest of $r_{i}$ is the Cartesian Forest $F(x[r_{i}+1\ldots r_{i+1}-1])$.
Figure 2: The Cartesian Forest associated to an ordered sequence $x$. Roots $r_{2}$ to $r_{4}$ are created since the sequence contains a value equal to the one at position 55. The values between two roots $r_{i}$ and $r_{i+1}$ belong to the tree enrooted in $r_{i}$. Therefore, except for the leftmost tree of a Cartesian Forest, a tree cannot have a left sub-forest. This idea is true at every level and can be seen in this example at positions $4$ and $12$.
4 Exact Cartesian Forest Matching
We adapt the linear representations of Cartesian Trees, such as the Parent-Distance [19] and the Skipped-Number representation [6] to Cartesian Forests. Basically, for a sequence $x$, when computing $C_{i}(x)$ from $C_{i-1}(x)$, the Parent-Distance of $i$ is equal to the distance between $i$ and its parent in $C_{i}(x)$ and its Skipped-Number is the number of nodes removed from the right branch of $C_{i-1}(x)$ comparing to the right branch of $C_{i}(x)$ (see [3]).
The main idea of the adaptation to Cartesian Forests is simple: both representations involve the comparisons of two values $x[j]$ and $x[k]$ in the original sequence $x$, for any $j<k$. Suppose we have either $x[j]>x[k]$ or $x[j]<x[k]$, then we are in the case of Cartesian Trees and the linear representation is the same. The idea for the Parent-Distance of the Cartesian Forest of a sequence $x$ is that its absolute value at a position $i$ gives the distance between a node $i$ and its parent or left sibling in the Cartesian Forest associated to the sequence $x[1\ldots i]$. Let $\operatorname{equal}_{x}(i)=\max(\{k\mid x[k]=x[i]\mbox{ with }1\leq k<i\}\cup\{0\})$.
Definition 6(Forest Parent-Distance representation $\operatorname{F-PD}_{x}$)
Given a sequence $x$ of length $m$, the Forest Parent-Distance representation of $x$ is an integer sequence $\operatorname{F-PD}_{x}[1\ldots m]$ defined as follows:
$\operatorname{F-PD}_{x}[i]=\begin{cases}i-\operatorname{small}_{x}(i)&\text{if }\operatorname{small}_{x}(i)>\operatorname{equal}_{x}(i)\cr-(i-\operatorname{equal}_{x}(i))&\text{if }\operatorname{small}_{x}(i)<\operatorname{equal}_{x}(i)\cr 0&\text{otherwise.}\end{cases}$
The referent table that we define next contains the position of the next smaller (or equal) value in the sequence.
Definition 7(The referent table $\operatorname{F-ref}_{x}$)
Given a sequence $x[1\ldots m]$, the referent table of $x$ is a sequence of sets $\operatorname{F-ref}_{x}[1\ldots m]$ such that
$\operatorname{F-ref}_{x}[i]=\begin{cases}\min_{i<j\leq m}\{j\mid x[j]\leq x[i]\}&\text{if such }j\text{ exists}\cr-1&\text{ otherwise.}\end{cases}$
Definition 8(Forest Skipped-Number representation $\operatorname{F-SN}_{x}$)
Given a sequence $x[1\ldots m]$, the Forest Skipped-Number representation of $x$ is an integer sequence $\operatorname{F-SN}_{x}[1\ldots m]$ such that
$\operatorname{F-SN}_{x}[i]=\begin{cases}|\{j<i\mid\operatorname{F-ref}_{x}[j]=i\}|&\mbox{if }\operatorname{small}_{x}(i)>\operatorname{equal}_{x}(i)\cr-|\{j<i\mid\operatorname{F-ref}_{x}[j]=i\}|&\mbox{if }\operatorname{small}_{x}(i)<\operatorname{equal}_{x}(i)\cr\end{cases}$
Figure 3 shows examples of Forest Parent-Distance representations, Forest Skipped-Number representations and referent tables of two Cartesian Forests.
Figure 3: Two sequences $x$ and $y$, their associated Cartesian Forests $F(x)$ and $F(y)$, and their corresponding Forest Parent-Distance representations, Forest Skipped-Number representations and referent tables. As one can see, $x$ is a prefix of $y$ and the forest $F(x)$ is transformed into a sequence of left subtrees in $F(y)$.
Given two sequences $x$ and $y$, we will denote $x\approx_{CF}y$ when the two sequences share the same Cartesian Forest.
Definition 9(Cartesian Forest Matching (CFM))
Given two sequences $p[1\ldots m]$ and $t[1\ldots n]$, find every position $j$, with $1\leq j\leq n-m+1$, such that $t[j\ldots j+m-1]\approx_{CF}p[1\ldots m]$.
Example 1
Let $t=(5,7,3,6,3,7,2,8,2,4,3,3)$ and $p=(2,3,1,4,1,5)$ respectively be the text and the pattern. We have two occurrences of $p$ in $t$ as $t[1\ldots 6]=(5,7,3,6,3,7)\approx_{CF}p$ and $t[5\ldots 10]=(3,7,2,8,2,4)\approx_{CF}p.$
Proposition 1
Algorithm 1 solves the CFM problem in $\mathcal{O}(n)$ space and $\mathcal{O}(mn)$ time in the worst-case and $\mathcal{O}(n)$ time on average.
Proposition 2
Using a Cartesian Forest border table, the CFM problem can be solved in $\mathcal{O}(n)$ time and space by adapting the KMP approach used in [19].
Note that, in order to prove both propositions, it is sufficient to prove that two sequences share the same Cartesian Forest if and only if they share the same linear representation. It can done by adapting the proof from [19]111see Appendix A.1.
5 Approximate Cartesian Forest Matching
The above approach for exact CFM can be extended to approximate CFM with one difference in a similar way as approximate CTM with one difference is done in [3]. Given a sequence $x$ of length $m$ and $i$ a position $1\leq i<m$, let $y$ be the sequence defined by swapping the elements in positions $i$ and $i+1$ in $x$. In [3], it is shown that there are at most $3$ mismatches between the Skipped-number representation of the Cartesian Tree of $x$ and the Skipped-number representation of the Cartesian Tree of $y$. The same result holds for the Forest Skipped-Number representation of the Cartesian Forest of $x$ and the Forest Skipped-Number representation of the Cartesian Forest of $y$ considering that two elements match if their absolute values are equal. Thus, based on the results in [3], the CFM with one swap of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
Approximate CTM with one mismatch, one insertion or one deletion is solved by comparing the Parent-Distance representations from left to right and the Parent-Distance representations from right to left of the Cartesian Trees of $x$ and $y$. The Forest Parent-Distance representation from right to left of a Cartesian Forest can be defined in a similar way as for a Cartesian Tree. And then, approximate CFM with one mismatch, one insertion or one deletion can be solved in a similar way than approximate CTM with one mismatch, one insertion or one deletion. Thus, based on the results in [3], the CFM with one mismatch, one insertion or one deletion of a pattern of length $m$ in a text of length $n$ can be done in time $O(mn)$ in the worst case and in linear time in average.
6 Combinatorics of the Cartesian Forests
6.1 Recursive Definition
Definition 10(Cartesian Forest $F$)
A Cartesian Forest $F$ can be:
  • empty,
  • a sequence of $k$ trees rooted in $(r_{1},\ldots,r_{k})$, with $k\geq 1$, such that $r_{1}$ has a left Cartesian sub-Forest and a right Cartesian sub-Forest, and $r_{i}$ has only a right Cartesian sub-Forest for all $i\in\{2,\ldots,k\}$ (the left sub-Forest is necessarily empty).
Figure 4 shows examples of one Cartesian Forest and two forests that are not Cartesian.
Figure 4: On the left is a valid Cartesian Forest. In the middle, it is not a valid Cartesian Forest because the second tree in the right sub-forest of $r^{\prime}_{1}$ has a left sub-forest. On the right, the forest is not Cartesian: the second tree has a left sub-forest.
[…]
6.2 Generating Function
Let $F(z)=\sum_{n\geq 0}f_{n}z^{n}$ be the generating function of Cartesian Forests, where $f_{n}$ counts the number of Cartesian Forests with $n$ nodes, and $S(z)$ be the generating function of siblings. The translation between the recursive decomposition is made using the Symbolic Method [8]:
$\begin{cases}F(z)=z\cdot F^{2}(z)\cdot S(z)+1\\ S(z)=z\cdot F(z)\cdot S(z)+1\end{cases}$
from which we obtain that $S(z)=\frac{1}{1-z\cdot F(z)}$ and $F(z)=1+\frac{z\cdot F^{2}(z)}{1-z\cdot F(z)}$.
It can be shown from here that $F(z)=\frac{1+z+\sqrt{1-6z+z^{2}}}{4z}$, whose associated coefficients are known to be Schröder–Hipparchus numbers, also called super-Catalan numbers (Sequence $A001003$ on OEIS) enumerated by the following formula: $f_{n}=\sum_{i=1}^{n}\frac{1}{n}\binom{n}{i}\binom{n}{i-1}2^{i-1}$, which, amongst other things, counts the number of ways of inserting parentheses into a sequence of $n+1$ symbols, where each pair of parentheses surrounds at least two symbols or parenthesized groups, and without any parentheses surrounding the entire sequence. In [4], the authors show that:
$f_{n}\sim\frac{\sqrt{3\sqrt{2}-4}}{4\sqrt{n^{3}\pi}}(3+2\sqrt{2})^{n}\sim 0.07(5.828)^{n}n^{-\frac{3}{2}}.$
In the following subsections, we describe bijections between Cartesian Forests and classical combinatorial objects counted by the Schröder-Hipparchus numbers. We recall that since those objects all share the same generating function, there exists a bijection between those sets [^2] Figure 8 (Appendix) shows the one to one correspondence for $n\in\{1,2,3\}$ between Cartesian Trees, Schröder Trees and Parentheses Words. Any injection from one set to the other is a bijection.
6.3 A bijection with Schröder Trees
Definition 11(Schröder Tree $ST$)
A Schröder Tree is a tree whose internal nodes have two or more subtrees.
Schröder Trees with $n+1$ leaves are counted by $f_{n}$. Given a Cartesian Forest $F$ with $k$ roots $(r_{1},\ldots,r_{k})$, we denote $\operatorname{left}(r_{1})$ the left sub-forest of the root $r_{1}$. For $i\in\{1,\ldots,k\}$, we denote $\operatorname{right}(r_{i})$ the right sub-forest of $r_{i}$. From a Cartesian Forest with $n\geq 1$ nodes, Algorithm 2 gives a recursive approach to building a Schröder Tree with $n+1$ leaves.
Input : A Cartesian Forest $F$ w. $n$ nodes and $k$ roots, s.t. $F=(r_{1},\ldots,r_{k})$
Output : A Schröder Treewith $n+1$ leaves
1 $ST\leftarrow$ a root;
2 if $F$ is not empty then
3 $(c_{1},\ldots,c_{k+1})\leftarrow$ Create a tuple of subtrees;
[…]
5 for $i\in\{2,\ldots,k+1\}$ do
[…]
7
8 $ST\leftarrow$ Add subtrees $(c_{1},\ldots,c_{k+1})$ to $ST$;
9
10return STST;
Algorithm 2 CFtoST(F)(F)
Input : A Cartesian Forest $F$ with $n$ nodes and $k$ roots, s.t. $F=(r_{1},\ldots,r_{k})$
Output : A sequence with parentheses and n+1n+1\ $\square$
1 $w\leftarrow\square$;
2 if $F$ is not empty then
3 $(w_{1}\cdots w_{k+1})\leftarrow$ Create a tuple of words ;
4 $w_{1}\leftarrow\textsc{CFtoW}(\operatorname{left}(r_{1}))$;
5 for $i\in\{2,\ldots,k+1\}$ do
6    $w_{i}\leftarrow\textsc{CFtoW}(\operatorname{right}(r_{i-1}))$;
7
8 $w\leftarrow\bm{(}w_{1}\cdot w_{2}\cdots w_{k+1}\bm{)}$;
9
10return w;
Algorithm 3 CFtoW($F$)
Figure 5: Algorithms that take a Cartesian Forest as input and return the corresponding Schröder Tree (algorithm on the left) or Parentheses Word (algorithm on the right). As one can see, both methods are very similar. In both cases, two different inputs imply two different outputs. Hence both functions are injections. Since there is the same number of objects of the same size in all three cases, the functions are bijections.
Algorithm 2 is an injective function: at each level, the number of subtrees added to the Schröder Tree is exactly $k+1$, where $k$ is the number of trees in the Cartesian Forest. Since both objects share the same generating function, we obtain the result announced in Lemma 1.
6.4 A bijection with Parentheses Words
Definition 12(Parentheses Word $w$)
[…]
Do note that, in this definition, unlike the more commonly found definition of these words, we allow parentheses to surround the entire sequence. We notably do so in order to simplify Algorithm 3 and to make the bijection more apparent to the reader. But since every parentheses word that is not $\square$ is contained between a pair of parentheses, it is not necessary to represent it. Informally, one may consider the symbols of the word as separators between the nodes and each group of parentheses as a (sub-)forest. Using the same arguments as for Schröder Tree, Algorithm 3 is an injective function.
Lemma 1
Algorithms 2 and 3 are bijective functions that map a Cartesian Forest with $n$ nodes to a Schröder Tree with $n+1$ leaves (Algorithm 2), and to a Parentheses Word with $n+1$ symbols (Algorithm 3). They both have a $\Theta(n)$ time-complexity and a $\Theta(n)$ worst-case space complexity.
7 Cartesian Forest Signature and Cartesian Forest Matching using a filter
In [6], the authors propose a signature (or perfect hash) of a Cartesian Tree, based on the Skipped-number representation. Given a Cartesian Tree with $n$ nodes, its signature is an integer with at most $2n$ bits. In this section, we extend this notion of signature to Cartesian Forests, obtaining a signature with at most $3n$ bits.
Definition 13(Cartesian Forest Signature)
Given the Forest Skipped-Number representation $\operatorname{F-SN}_{x}$ of a sequence $x$ (see Definition 8), its signature is defined in the following way: given a position $i\in\{1,\ldots,n\}$ in the Forest Skipped-Number representation, each value $\operatorname{F-SN}_{x}[i]$ is encoded by a sequence of bits, that are concatenated to obtain an integer.
  • The first bits concern the sign of $\operatorname{F-SN}_{x}[i]$: If $\operatorname{F-SN}_{x}[i]=0$, it is equal to $\tt 0$}; The case $\operatorname{F-SN}_{x}[i]<0$ is encoded by $\tt 10$}, and the case $\operatorname{F-SN}_{x}[i]>0$ is encoded by $\tt 11$}.
  • If $\operatorname{F-SN}_{x}[i]\neq 0$, the following bits are a unary encoding of $|\operatorname{F-SN}_{x}[i]|$: $|\operatorname{F-SN}_{x}[i]|-1$ bits equal to $\tt 1$} followed by a bit equals to $\tt 0$}.
Since the total number of skipped nodes cannot exceed $n$, a signature contains at most $3n$ bits. This representation could be used efficiently as a perfect hash for small patterns, to obtain an efficient algorithm to solve the Cartesian Forest Matching problem: one only needs to update the signature as one would update the Forest Skipped-Number representation of the chunk of text compared to the pattern, using bitwise operations. But it becomes inefficient when $3n$ exceeds the size of a register, as one cannot use those operations efficiently anymore. Therefore, in order to accelerate the Cartesian Forest Matching in a Rabin-Karp fashion we now introduce a filtering method.
Definition 14(Cartesian Forest $\tau$-Filter)
Given a sequence $x$ of length $n$ and its linear representation $\operatorname{LN}_{x}$, its $\tau$-Filter, denoted by $\mathcal{F}il_{x}$, is a sequence of $\tau$ bits such that $\mathcal{F}il_{x}[i]=0$ if $\operatorname{LN}_{x}[n-\tau+i]=0$ and $\mathcal{F}il_{x}[i]=1$ otherwise.
Algorithm 1 can be adapted: line $5$ is only tested if $\mathcal{F}il_{x}$ is equal to $\mathcal{F}il_{p}$. $\tau$ can be adapted to match the size of a classical register, that is $32$, $64$ or $128$ bits. Therefore, comparing two filters can be made in constant time. The update of the filter can also be made in constant time by tracking the positions in $\operatorname{LN}_{x}$ that have been updated (see [3] for more details on the update function of the linear representations). In the following Section, we only implemented the filter for the Forest Skipped-Number representation.
8 Experiments
In this Section, we implemented a Cartesian Forest $\tau$-Filter using $\tau=64$. The first set of experiments considers the uniform distribution over text and patterns of respective length $n$ and $m$, over a $k$ letter alphabet. Figure 6 sums up the details and observations. As one can see, the Forest Parent-Distance version is the slowest (PD meta), then comes the one using the Forest Skipped-Number representation (SN meta). The method using a filter (SN hash) is even faster, but only with a slight margin. Finally, the KMP version [19] is the fastest one (PD KMP). Note that we tried to implement the Skip-Search algorithm [21] without the SIMD instructions, but the experimental results were disappointing. The average cost slightly increases with $k$, which is probably due to the cost of the update function, since the average Forest Parent-Distance increases with the size of the alphabet. Note that, whatever the linear representation is, the number of comparisons at Line $5$ in Algorithm 1 is the same, since they encode the same prefix. Therefore, it is likely that the extra cost of the Forest Parent-Distance representation is due to the fact that both requires to use a stack in their implementations but more values are pushed in the stack, in the Forest Parent-Distance case.
Figure 6: We randomly generated 10 00010\,000 patterns of size $m$ and texts of size 1 0001\,000. The $x$-axis represents $m$ and the $y$-axis is the average time taken by each algorithm in microseconds. From left to right, the alpĥabet size is respectively equal to 22, 44 and $m$.
The experiment in Figure 7 uses the random generator from [5]. It shows the algorithm remains efficient even with (not too) low entropy. It generates sequences uniformly amongst those of a fixed length over a fixed alphabet and a fixed Collision Entropy (that is Rényi Entropy with $\alpha=2$). The Collision Entropy is a function of the probability for two random variables (a letter in the pattern $p$ and the text $t$) to be equal. The lower the entropy is, the higher average complexity of the algorithms. As a matter of fact, if the entropy reaches 0, then both the text and the pattern contains the repetition of a unique symbol, which corresponds to the worst-case complexity of Algorithm 1. Though, as one can see in Figure 7, the average cost of the algorithms quickly drops and stabilizes. The KMP version is the most efficient (this is in line with the results of [2] showing that the KMP algorithm is efficient for order preserving matching) and is (unsurprisingly) very stable, even when the entropy is low.
Figure 7: For each value of the collision entropy ($x$-axis), 10 00010\,000 random texts of length 1 0001\,000 and patterns of length $100$ were generated, and the four algorithms were applied to them. On the left $m=10$, and on the right $m=100$. The $y$-axis is the average time in microseconds. The added value of using the filter method, compared to the simple Forest Skipped-Number version, is more important when the entropy is low, whereas the difference in efficiency between the Forest Parent-Distance version and the Forest Skipped-Number version drops.
Appendix 0.A Appendix
0.A.1 On the Linear Representations of Cartesian Forests
We show that two sequences $x$ and $y$ share the same Cartesian Forest if and only if they share the same Forest Parent-Distance representation.
Proof
If the sequences have different lengths, then they have different Cartesian Forests and different Forest Parent-Distance representations and it holds. We need only consider the case when $x$ and $y$ have the same length $m$. We use a proof by induction on $m$.
For $m=0$, both Cartesian Forests and their representations are empty and trivially match. For $m=1$, both Cartesian Forests have only one root and are the same, and their Forest Parent-Distance representation are both equal to 0. The property holds for $m=1$. Let us now assume that the property holds for a certain $m=k$ and move onto $m=k+1$.
[…]
  • If positions $k+1$ have a parent in both Cartesian Forests, let $\alpha$ (resp. $\beta$) be the parent of $k+1$ in $F(x[1\ldots k+1])$ (resp. $F(y[1\ldots k+1])$). Since both sequences share the same Cartesian Forest, we must have $\alpha=\beta$. We have $\operatorname{F-PD}_{x}[1\ldots k]=\operatorname{F-PD}_{y}[1\ldots k]$ by induction hypothesis. If $\alpha$ and $\beta$ are parents then $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]=k+1-\alpha=k+1-\beta$. Therefore $\operatorname{F-PD}_{x}[1\ldots k+1]=\operatorname{F-PD}_{y}[1\ldots k+1]$.
  • If positions $k+1$ have siblings in both Cartesian Forests, let $\alpha$ (resp. $\beta$) be the sibling of $k+1$ in $F(x[1\ldots k+1])$ (resp. $F(y[1\ldots k+1])$). The same reasoning as before applies, except $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]=-(k+1-\alpha)=-(k+1-\beta)$. Therefore $\operatorname{F-PD}_{x}[1\ldots k+1]=\operatorname{F-PD}_{y}[1\ldots k+1]$.
  • If positions $k+1$ possess neither parents or siblings, then we have $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]=0$ and thus $\operatorname{F-PD}_{x}[1\ldots k+1]=\operatorname{F-PD}_{y}[1\ldots k+1]$.
[…]
  • If $\operatorname{F-PD}_{x}[k+1]=\operatorname{F-PD}_{y}[k+1]\neq 0$, let $\alpha=k+1-|\operatorname{F-PD}_{x}[k+1]|$.
    • If $\operatorname{F-PD}_{x}[k+1]<0$, then $\alpha$ is the sibling of $k+1$ in both $F(x[1\ldots k+1])$ and $F(y[1\ldots k+1])$.
    • Otherwise, if $\operatorname{F-PD}_{x}[k+1]>0$, that means $\alpha$ is the parent of $k+1$ in $F(x[1\ldots k+1])$ and $F(y[1\ldots k+1])$. Also the right sub-Forest of $\alpha$ becomes the left sub-Forest of $k+1$.
  • If $\operatorname{F-PD}_{x}[k+1]=0$, $k+1$ is the root of $F(x[1\ldots k+1])$ and $F(x[1\ldots k])$ becomes the left sub-Forest of $x[k+1]$. We derive $F(y[1\ldots k+1])$ from $F(y[1\ldots k])$ in the same way.
In all three cases, we can conclude that $F(x[1\ldots k+1])=F(y[1\ldots k+1])$.
Hence, there is a one-to-one mapping between Cartesian Forests and Forest Parent-Distance representations.
A similar proof using the same arguments can be used to prove a one-to-one mapping between Cartesian Forests and the Forest Skipped-Number representation.
0.A.2 On the Symbolic Method
The Symbolic Method was introduced in [8]. It is a method to directly translate a recursive definition of a combinatorial class into a generation function. Recall that a combinatorial class is a set of objects associated with a size function, such that for all $n\geq 0$, the subset of objects of size $n$ is finite. This section only contains the information necessary to understand how we obtained the generating function $F(z)$. Therefore, we focus on unlabelled objects and simple operations. A reader willing to learn more should read [8].
Let $\mathcal{C}$ be a combinatorial class and $\mathcal{C}_{n}$ be the set of objects of $\mathcal{C}$ of size $n$. We have $\mathcal{C}=\bigcup_{n\geq 0}\mathcal{C}_{n}$ The generating function $C(z)$ of $\mathcal{C}$ is defined as a power series
$C(z)=\sum_{n\geq 0}c_{n}z^{n},$
where $c_{n}=|\mathcal{C}_{n}|$.
The symbolic method is a dictionary that translates operations on sets into operations onto their associated generating functions.
The Null size element.
Assume $\mathcal{C}=\mathcal{C}_{0}$, that is it only contains an object of size 0. Then $C(z)=1$.
The size 1 element.
Assume $\mathcal{C}=\mathcal{C}_{1}$, that is it only contains an object of size $1$. In our case, this would be a tree containing a unique node. Then $C(z)=z$.
Disjoint Union.
Assume $\mathcal{C}=D\cup E$, where $D$ and $E$ are two combinatorial classes. Then $C(z)=D(z)+E(z)$.
Cartesian Product
Assume $\mathcal{C}=D\times E$, where $D$ and $E$ are two combinatorial classes. Then $C(z)=D(z)\cdot E(z)$.
We recall the recursive decomposition that we obtained for Cartesian Forest:
[…]
The first combinatorial class $F$ is defined as:
  • either an empty forest (marked with $\emptyset$ whose associated generating function is $1$).
  • or a Cartesian product between a sequence of siblings, whose generating function will be denoted $S(z)$ and a tree, that is decomposed into three parts:
    • a root, that is a tree of size $1$, whose generating function is $z$
    • a left and a right sub-forest, that is to say two Cartesian Forests.
    The generating function of this tree is therefore equal to $F(z)\cdot z\cdot F(z)$.
Using the Symbolic Method we therefore obtain that $F(z)=1+z\cdot F^{2}(z)\cdot S(z)$. Using the same ideas, we obtain that
$S(z)=z\cdot F(z)\cdot S(z)+1$
$\implies S(z)-z\cdot F(z)\cdot S(z)=1$
$\implies S(z)=\frac{1}{1-z\cdot F(z)}$
And therefore
$F(z)=1+\frac{z\cdot F^{2}(z)}{1-z\cdot F(z)}$
$\implies F(z)\cdot(1-z\cdot F(z))=1-z\cdot F(z)+z\cdot F^{2}(z)$
$\implies F(z)-z\cdot F^{2}(z)=1-z\cdot F(z)+z\cdot F^{2}(z)$
$\implies 2z\cdot F^{2}(z)-(1+z)\cdot F(z)+1=0$
Solving this equation we obtain $\Delta=b^{2}-4ac=(1+z)^{2}-8z=1-6z+z^{2}$ and
$F(z)=\frac{-b+\sqrt{\Delta}}{2a}=\frac{1+z+\sqrt{1-6z+z^{2}}}{4z}$
Note here that we do not consider $\frac{-b-\sqrt{\Delta}}{2a}$ because we know $F(z)$ must have positive coefficients.
0.A.3 Bijection with Schröder Trees and Parentheses Words
[…]
Figure 8: Correspondence between Cartesian Forests with $n$ nodes, Schröder Trees with $n+1$ leaves and sequences of length $n+1$ with parentheses when $n\in\{1,2,3\}$ (respect. top, middle, bottom). $\square$ symbols represent leaves and $\circ$ symbols represent internal nodes.